OpenStudy (anonymous):

integral x^2 logx dx

OpenStudy (zarkon):

parts

OpenStudy (lgbasallote):

im thinking integration by parts... let u = logx dv = x^2 du = 1/x v = (x^3)/3 \(\Huge \frac{x^3 logx}{3} - \int{}{} \frac{x^3}{3x}\)

OpenStudy (lgbasallote):

can you fo the rest from there?

OpenStudy (anonymous):

yeah i think so but what is the integral of just logx for reference?

OpenStudy (zarkon):

\[\int\ln(x)dx=x\ln(x)-x+c\]

OpenStudy (anonymous):

does that apply for common logs too or just nat logs

OpenStudy (zarkon):

\[\log_a(x)=\frac{\ln(x)}{\ln(a)}\]

OpenStudy (zarkon):

\[\int\log_a(x)dx=\int\frac{\ln(x)}{\ln(a)}dx=\frac{1}{\ln(a)}\int\ln(x)dx\]

OpenStudy (anonymous):

thank you!